find an equation of the line containing the given pair of points ( 4,2) (12,8) y= I came up with y= 4 am I close
Answer by stanbon(75887)

find an equation of the line containing the given pair of points ( 4,2) (12,8)
-------------
No, you are not close.
--------------
slope = (8-2)/(12-4) = 6/8 = 3/4
-----------------------------------
Form is y = mx+b and you have m=3/4 and y = 2 when x = 4
---
2 = (3/4)*4 + b
b = -1
-----------------
Equation:
y = (3/4)x - 1
======================
Cheers,
Stan H.